LBC/LUCI Technical Observing, prep for Partner Science:
UT 2010 Jan 30
Telescope Operator: DHuerta
Instrument Support: JMorris
00:15 Shutdown and restart: LUCI's TelescopeService (build switch to BP7) and InstrumentManager (to correct lamp indicator problem). Also restarted GEIRS Server, but it probably did not need this.
LUCI1 functional checkout:
LUCI computer agrees with linux boxes
showSupervisorStatus shows all OK
Web-IO shows power on to MCE/ROE and Calib unit.
UI Startup OK, initialize instrument, image saved went to right places
Cannot check darks, dome lights on and we are opening.
Lamp indicators now come on.
0004-0033 K 6k @ start
0034-0035 H aborted, low counts
0036-0045 H 5.0s 3.8k @ start
0046-0055 J 5.0s 3.4k @ start
0056-0065 H 5.0s 2.0k @ start
0066-0075 J 5.0s .k @ start
0076 ptcorr, collimating. Started with some coma and it took oddly long to converge. Is this something that the new wfs algorithm will improve?
Repeat check of dithering along slit (20x 2' dither)
slitdither.acq at BS9108. slitdither.sci. Lost guide star at dither position #14. Previous dithers were OK. What happened?
started at the pointing model: IE -115.7831 CA +20.4825
and offset to : IE -90.01550 CA -20.30070
Why are we so far off the pointing model made just a few days ago?
Restarted slitdither. Acquisition took 7 minutes, of which 3 minutes were integration and 1.5 minutes was moving the mask to the FPU from turnout.
Test rangebal script...worked. Why are we so far off the pointing model?
Old Delta New
------- ------- -------
IE -90.016 18.069 -71.946
CA -20.301 2.852 -17.448
Astrometric fields: Ks 5x10s
BS9108 0118, 0119
BS9116 0120, 0121
BS9118 0122, 0123
03:43 BS9116 Collect through-focus sequence
Andrew says to double the offsets, and I'll scale back to 30s exposures to go a little faster...
Send ACTIVE preset to the off-axis guide star here and check how much Z4 is applied...almost nothing.
ACTIVE preset on-axis and allow to converge. Mirror "integrator" not behaving well. Collimated on-axis, image with the guide probe in the way is #136. Then guide-mode preset to off-axis guide star is #137. Use earlier sky frame this field.
First active collimation after this sent 500nm. Image taken after converged is #143.
04:56 Field aberration data collection on BS9126, EL~73 and rising.
05:55 Restarted again at #8 after an ACTIVE preset from LUCI and allow it to collimate on axis (last wfs 192nm). Problem with the preset (GCS returned a guidebox-sized image for acquisition). Recollimating. Restarted 06:07.
07:29 Started field_collect on BS9118. Doug corrected his script to read a better measure of the collimation state of the telescope and it seems to be working correctly. Seeing started out good but rose to ~1.3" by the end of the script.
We are seeing drifts in the telescope pointing even at low EL. It is mostly in CA, but quite large. We are currently ~75 arcsec off the pointing model. This is troubling because we only made this pointing model a few nights ago and it had 2" rms.
09:02 M82.acq after pointing correction on ACT0404 nearby. Check exposure time for galaxy...
- test expo 1x10s #157. saturated in core
- test expo 1x4s #158 Use 15x4.0s
- M82sky 159-161
- M82ctr 162-166
- M82sky 167-169 TRACK mode, #168 has a jump, #169 jumped back!
- M82ne 170-174
- M82sky 175-177
- M82sw 178-182
- M82sky 183-185
- M82ctr 186-190
- M82sky 191-193
10:30 BS9145 to run science-type dice5 pattern
- Execute script: ~/LBTO/SciencePrep/dice5_10x (~1 hr)
- Calculate rms/position, scale & rotation of pattern
10:31 Note: the first preset to BS9145 completed, but was then cancelled shortly thereafter because "telescope off position for too long". Was this another jump?
10:37 imagepair (for astrometry and to check field in advance of dice5) #193, 194
10:40 dice5_10x at BS9145 Lost guiding (again) at image 25/50. DGH will submit a critical InformationTechnology
as we have seen this 5-6 times tonight.
Restarted dice5_10x script.
11:39 Saw on guider another jump. The mount and guider were out of phase for a bit there, but it recovered without cancelling the preset. See LUCI#0228.
11:51 Offset failed at image 19/25 because guide loop did not pause within the timeout. Seeing getting worse.
12:08 Some more astrometric images:
- BS9148 240, 241
- BS9152 242 243
- BS9156 244, 245
- BS9158 246, 247
- BS9162 251, 252, 253, 254
13:15 Test acquisition for Matthias Mueller. EL~42 Seeing is bad, jumpy. Images 260 - 263 (more?)
LUCI Observing Plan
task and verify it does what is needed on LUCI.
Additional work on Andrew's M1/M2 collimation?
Take a set of (monocular) range-balanced collimation data
- Stable conditions!
- implement this collimation model
Photometric zero point and extinction monitoring
Collect a "movie" of 2s LUCIFER images while an in-field offset is taking place.
- Check to make sure offsets still happen in a reasonable manner
Collect astrometric field data to check AGw calibration for monitoring
Improve map of AGw patrol field limits at high field angle
- Requires set up work on a stone field
Gather NB illumination correction data (BrG done on 20100606)
Re-test updatePointingReference functionality.
Continue gaining experience with LUCIFER on-sky (e.g. spec acq.)
Test dithering between extreme slits on MOS mask + telluric
- Manual needs 0.59836mm/" scale for the dXdY calculation
- The how-to set up telluric observations needs updating for clarity.
Investigate variable imaging FF gradients by taking data while the telescope is offset with respect to the enclosure (will need manual overriding of normal operations) to see if the observed gradients could be caused by something 'external to the instrument'.
- 28 Jan 2011